Solve for x and y.
2x+3y = 6
5x+2y = 15

Answer by Alan3354(69443)

Solve for x and y.
2x+3y = 6 times 2 --> 4x + 6y = 12
5x+2y = 15 times 3 -->15x + 6y = 45
--------------
4x + 6y = 12
15x + 6y = 45
-------------------- Subtract